If you have an average of 85. You have only 4 of the 5 grades that is figured into the grade, how do you find the 5th grade? The 4 grades are 78, 82, 94, and 95.
Answer by ewatrrr(24785)


Hi
Let x represent the 5th grade
(349 + x)/5 = 85
349+x = 425
x = 76, the 5th grade
CHECKING our Answer***
(78+ 82+94+95+76)5 = 85